# Munich, Germany PT43D

### v1

@@ 2 Related Work
Probabilistic 3D Shape Generation. More recently, researchers have explored the use of generative neural networks [Ho et al.(2020)Ho, Jain, and Abbeel, Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer, Goodfellow et al.(2020)Goodfellow, Pouget-Abadie, Mirza, Xu, Warde-Farley, Ozair, Courville, and Bengio, Van Den Oord et al.(2016b)Van Den Oord, Kalchbrenner, and Kavukcuoglu, Van Den Oord et al.(2016a)Van Den Oord, Dieleman, Zen, Simonyan, Vinyals, Graves, Kalchbrenner, Senior, Kavukcuoglu, et al.] to model the probabilistic distribution of 3D shapes conditioned on a single image, enabling the sampling of multiple hypotheses from the learned distribution. For instance, AutoSDF [Mittal et al.(2022)Mittal, Cheng, Singh, and Tulsiani] models shape distribution via an autoregressive process and combines it with a separately-trained image-conditioning model to generate shapes based on single RGB images. SDFusion [Cheng et al.(2023)Cheng, Lee, Tulyakov, Schwing, and Gui] learns to sample from a target shape distribution by reversing a progressive noise diffusion process, generating single-view reconstructions via classifier-free image guidance [Ho and Salimans(2022)]. These approaches typically focus on images with fairly clear and complete visual descriptions of the object, but unfortunately do not address common real-world scenarios where objects are captured under strong occlusion or truncation. Our method, instead, takes advantage of generative neural networks, analyzing common realistic scenarios where images are highly occluded or truncated.
@@ 3.1 Method Overview
Given an input image $c$, an empty query sequence $s_{q}$ in the length of $n+1$ and a start token $<sos>$, our model generates the shape distribution over the latent space $E$ by filling the query sequence autoregressively. The complete sequence $s_{c}$ comprises of discrete grids, where each location $i$ contains the distribution $p_{i}(e)$ of 3D latent features. During inference, we can sample from $p_{i}(e)$ to acquire shape feature index $q_{i}$ for each location and further get the corresponding shape feature $e_{i}$ from the codebook $Z$. Finally, a pretrained decoder $D_{\phi}$ utilizes these shape features to produce the reconstruction output. The overall framework, as displayed in Fig 2, comprises three key components: the image encoder, the conditional cross-attention module and the autoregressive modeling with transformer.
Figure 2: Method Overview. Our model adopts a low-dimensional and discretized latent space for shape generation. Given an input RGB image, it is first processed by an image encoder. Then the extracted image encodings are fed into the conditional module where they attend with the input sequence consisting of a start token, potentially filled grids and a target grid via cross-attention such that the output sequence contains essential image features. Then the transformer processes this sequence and predicts the latent feature for the target grid. Conditional cross-attention and transformer are adopted autoregressively until a complete sequence is output where each cell contains the conditional distribution of latent features. During inference, we can gather the complete sequence into grids then sample and decode to multiple plausible hypotheses.
@@
Figure 3: Identifying Most Relevant Regions in the Image via Conditional Cross-Attention. We perform cross-attention between image encodings and the input sequence to effectively identify the most relevant region of interest from the input image. The input sequence is multiplied with a weight matrix $Q$ and the image encodings are multiplied with weight matrices $V$ and $K$ independently. The resulting sequence contains a weighted sum of image features. We then project it to the original dimension and concatenate with the original sequence embeddings.
Conditional Cross-Attention Module. The conditional module establishes a connection between the image encodings and the grid sequence $s_{t}$ at current autoregressive step $t$, ensuring that the output grid sequence incorporates the essential image features for subsequent operations. Fig. 3 depicts the architecture of it. $s_{t}$ comprises of the start token $<sos>$, potentially filled grids and the target grid with its position embedding. We perform cross-attention between $s_{t}$ and the image encodings, where each grid in $s_{t}$ serves as a query. The output sequence after cross-attention is essentially a weighted sum of image features. Then we project it to the original dimension and concatenate it with original embeddings.
Autoregressive Modeling with Transformer. After obtaining the sequence that contains the necessary image features at step $t$, we feed it into the transformer and predict the feature index $q_{l_{t}}$ for current target location $l_{t}$. We learn this model by maximizing the log-likelihood of the latent representations and perform this process (including conditional cross-attention step) autoregressively until all locations are filled.
@@ 3.2 Training and Inference
$p_{l_{t}}(e)=T_{\theta}(\left\{\hat{e_{l_{t-1}}},\hat{e_{l_{t-2}}},\hat{e_{l_{t-3}}},\cdots\right\},l_{t},c).$ (1)
where $l_{t}$ represents current location and $\hat{e_{l_{t}}}$ is generated by sampling from the distribution $p_{l_{t}}(e)$. We then feed the complete sequence into $D_{\phi}$ to generate one plausible 3D shape.
@@ 5.3 Ablation Studies
6 Conclusion and Future Work
In this work, we propose a new approach for generating the probabilistic distribution of 3D shape reconstructions conditioned on a highly ambiguous RGB image. We employ cross-attention to effectively identify the most relevant region of interest from the input image for shape generation. We also introduce a synthetic data augmentation approach to enable effective multi-hypothesis generation, by computing multiple potential ground-truth shapes that correspond to a single image observation during training. This synthetic augmentation enables pre-training that notably improves shape reconstruction results when fine-tuned on real-world data. Experiments demonstrate that our model outperforms state of the art in both generation quality and plausible generation diversity. For future work, we consider expanding the generation scale from object-level to scene-level, leveraging monocular views.

### v3

@@ 2 Related Work
Probabilistic 3D Shape Generation. More recently, researchers have explored the use of generative neural networks [Ho et al.(2020)Ho, Jain, and Abbeel, Rombach et al.(2022)Rombach, Blattmann, Lorenz, Esser, and Ommer, Goodfellow et al.(2020)Goodfellow, Pouget-Abadie, Mirza, Xu, Warde-Farley, Ozair, Courville, and Bengio, Van Den Oord et al.(2016b)Van Den Oord, Kalchbrenner, and Kavukcuoglu, Van Den Oord et al.(2016a)Van Den Oord, Dieleman, Zen, Simonyan, Vinyals, Graves, Kalchbrenner, Senior, Kavukcuoglu, et al.] to model the probabilistic distribution of 3D shapes, enabling the sampling of multiple hypotheses from the learned distribution. For instance,  [Yan et al.(2022)Yan, Lin, Mitra, Lischinski, Cohen-Or, and Huang, Ibing et al.(2021)Ibing, Kobsik, and Kobbelt, Siddiqui et al.(2024)Siddiqui, Alliegro, Artemov, Tommasi, Sirigatti, Rosov, Dai, and Nießner, Nash et al.(2020)Nash, Ganin, Eslami, and Battaglia, Zhang et al.(2022)Zhang, Nießner, and Wonka] utilize transformer-based architecture to generate shapes, achieving impressive results. With recent advancements in denoising diffusion models,  [Mo et al.(2024)Mo, Xie, Chu, Hong, Niessner, and Li, Hui et al.(2022)Hui, Li, Hu, and Fu, Gupta et al.(2023)Gupta, Xiong, Nie, Jones, and Oğuz, Zhang et al.(2023)Zhang, Tang, Niessner, and Wonka] generate high-quality shapes by iteratively refining noise-corrupted data through a reverse diffusion process guided by learned probability distributions. For image-conditioned shape generation, AutoSDF [Mittal et al.(2022)Mittal, Cheng, Singh, and Tulsiani] models shape distribution via an autoregressive process and combines it with a separately-trained image-conditioning model to generate shapes based on single RGB images. SDFusion [Cheng et al.(2023)Cheng, Lee, Tulyakov, Schwing, and Gui] learns to sample from a target shape distribution by reversing a progressive noise diffusion process, generating single-view reconstructions via classifier-free image guidance [Ho and Salimans(2022)]. These approaches typically focus on images with fairly clear and complete visual descriptions of the object, but unfortunately do not address common real-world scenarios where objects are captured under strong occlusion or truncation. Our method, instead, takes advantage of generative neural networks, analyzing common realistic scenarios where images are highly occluded or truncated.
@@ 3 Methodology
3.1 Shape Compression and Discretization
For each shape $X$, a shape encoder $E_{\phi}(\cdot)$ encodes its Truncated Signed Distance Field (T-SDF) into a low-dimensional 3D grid feature $E\in g^{3}\times D$, where $g$ represents the resolution of the 3D grid, and $D$ denotes the dimension of the 3D grid features. We establish a discrete latent space containing $K$ embedding vectors, each with a dimension of $D$. For the shape feature $e_{i}$ at grid position $i$, a vector quantization operation $VQ(\cdot)$ is employed to find its nearest neighbor $e^{\prime}_{i}$ from the $K$ embedding vectors in the discrete latent space. The feature input into the shape decoder $D_{\phi}(\cdot)$ is the 3D discrete latent feature grid $E^{\prime}\in g^{3}\times D$, comprising discrete features $e^{\prime}_{i}$ at grid locations i. The formula is:
$E=E_{\phi}(X),E^{\prime}=VQ(E),X^{\prime}=D_{\phi}(E^{\prime}).$ (1)
[…]
@@ 3.2 Method Overview
Given an input image $c$, an empty query sequence $s_{q}$ in the length of $n+1$ and a start token $<sos>$, our model generates the shape distribution over the latent space $E^{\prime}$ by filling the query sequence autoregressively. The complete sequence $s_{c}$ comprises of discrete grids, where each location $i$ contains the distribution $p_{i}(e^{\prime})$ of 3D latent features. During inference, we can sample from $p_{i}(e^{\prime})$ to acquire shape feature index $q_{i}$ for each location and further get the corresponding shape feature $e^{\prime}_{i}$ from the codebook $Z$. Finally, a pretrained decoder $D_{\phi}$ utilizes these shape features to produce the reconstruction output. The overall framework, as displayed in Fig 2, comprises three key components: the image encoder, the conditional cross-attention module and the autoregressive modeling with transformer.
Figure 2: Method Overview. Our model adopts a low-dimensional and discretized latent space for shape generation. Given an input RGB image, it is first processed by an image encoder. Then the extracted image encodings are fed into the conditional module where they attend with the input sequence consisting of a start token, potentially filled grids containing latent features, and a target grid containing its position embedding via cross-attention such that the output sequence contains essential image features. Then the transformer processes this sequence and predicts the latent feature for the target grid. Conditional cross-attention and transformer are adopted autoregressively until a complete sequence is output where each cell contains the conditional distribution of latent features. During inference, we can gather the complete sequence into grids then sample and decode to multiple plausible hypotheses.
@@
Figure 3: Identifying Most Relevant Regions in the Image via Conditional Cross-Attention. We perform cross-attention between image encodings and the input sequence to effectively identify the most relevant region of interest from the input image. The input sequence consisting of a start token, potentially filled locations with their own predicted latent features, and a target location with its position embedding, is multiplied with a weight matrix $Q$ and the image encodings are multiplied with weight matrices $V$ and $K$ independently. The resulting sequence contains a weighted sum of image features. We then project it to the original dimension and concatenate with the original sequence embeddings.
Conditional Cross-Attention Module. The conditional module establishes a connection between the image encodings and the grid sequence $s_{t}$ at current autoregressive step $t$, ensuring that the output grid sequence incorporates the essential image features for subsequent operations. Fig. 3 depicts the architecture of it. $s_{t}$ comprises of the start token $<sos>$, potentially filled grids containing latent features, and the target grid with its position embedding. We perform cross-attention between $s_{t}$ and the image encodings, where each grid in $s_{t}$ serves as a query. The output sequence after cross-attention is essentially a weighted sum of image features. Then we project it to the original dimension and concatenate it with original embeddings.
Autoregressive Modeling with Transformer. After obtaining the sequence that contains the necessary image features at step $t$, we feed it into the transformer and predict the feature index $\hat{q_{l_{t}}}$ for current target location $l_{t}$. We then fill current target location $l_{t}$ with the latent feature corresponding to the predicted feature index $\hat{q_{l_{t}}}$. We learn this model by maximizing the log-likelihood of the latent representations and perform this process (including conditional cross-attention step) autoregressively until all locations are filled.
@@ 3.3 Training and Inference
$p_{l_{t}}(e^{\prime})=T_{\theta}(\left\{\hat{e^{\prime}_{l_{t-1}}},\hat{e^{\prime}_{l_{t-2}}},\hat{e^{\prime}_{l_{t-3}}},\cdots\right\},l_{t},c).$ (2)
where $l_{t}$ represents current location and $\hat{e^{\prime}_{l_{t}}}$ is generated by sampling from the distribution $p_{l_{t}}(e^{\prime})$. We then feed the complete sequence into $D_{\phi}$ to generate one plausible 3D shape.
@@ 5.3 Ablation Studies
5.4 Limitations
In our image-based shape generation framework, we utilize patch-wise image encodings as conditional input, enabling robust global alignment between input object observations and output reconstructions. However, this may not fully capture the shape variance and ambiguities that lie in the local details, such as the wheels of the chair present in Fig. 5. A potential approach to overcome this limitation involves adopting pixel-wise image features that are more locally extracted. In addition, while our method achieves promising reconstruction results at object-level, generating scene-level reconstructions could provide richer perception for applications like robotics. In our future work, we consider expanding the generation scale from object-level to scene-level, leveraging monocular views.
6 Conclusion
In this work, we propose a new approach for generating the probabilistic distribution of 3D shape reconstructions conditioned on a highly ambiguous RGB image. We employ cross-attention to effectively identify the most relevant region of interest from the input image for shape generation. We also introduce a synthetic data augmentation approach to enable effective multi-hypothesis generation, by computing multiple potential ground-truth shapes that correspond to a single image observation during training. This synthetic augmentation enables pre-training that notably improves shape reconstruction results when fine-tuned on real-world data. Experiments demonstrate that our model outperforms state of the art in both generation quality and plausible generation diversity. We hope our work will draw increased attention to this important and realistic problem, highlighting the need for innovative solutions in 3D shape reconstruction from single ambiguous image data.
7 Acknowledgements
This work was supported by the ERC Starting Grant SpatialSem (101076253).